# Informativity and Identifiability for Identification of Networks of Dynamical Systems

### v1

@@ IV Informativity
Here we consider $\Pi$ to just be a function of the free entries of the matrix MM. [^1] When no ambiguity arises we will often just refer to $M$ even if we mean the free entries of $M$. The only property we need regarding these free entries when we investigate generic rank is that they are algebraically independent [^2] Algebraically independent is the formal way of saying that the entries are free. commuting indeterminates over the coefficient field. The coefficient field can be taken as the field of real numbers if no transfer functions are known. In case some transfer functions are known, we take the coefficient field as the field of rational functions of transfer functions. This is a field with the standard definition of addition and multiplication. We will later on see that the generic rank is equal to the rank for almost all $M$. Since $\Pi$ is a rational function in $z$, it follows that
@@ IV-A Graphical Condition
is a proper affine variety in the set of all $M$. Hence it holds that $\operatorname{rank}_{g}\Pi=\operatorname{rank}\Pi$ for almost all $M$. Then it is straightforward to see that Theorem 2 of that paper can also be generalized. We are therefore able to derive a graph-theoretic characterization of the generic rank of $\Pi$.
@@ IV-B Gröbner Basis Condition
In case some of the entries of $M$ are known, then one can instead of polynomials with coefficients in the field $\mathbb{R}$ consider polynomials with coefficients in the field of rational functions of transfer functions in the known entries of $M$. This will, in most cases, speed up the computation of the Gröbner basis.
@@ V Identifiability
$F(X)=R_{y}P^{-1}Q_{r}+S_{yr}$
Similarly as in [15], we consider the domain $V_{o}$ of $F$ to be the set of all complex-valued vectors $X$ of dimension $k$ for which the inverse of $P$ exists. This is a dense open subset of $\mathbb{C}^{k}$. We then define the set of admissible $G_{c}$ as $V_{c}=F(V_{o})$. If we take $V_{c}$ to be the co-domain of $F$, it follows that $F$ is a surjective function.
@@
From this definition it can be shown that if a network is generically locally identifiable, then it can be recovered up to a discrete ambiguity, i.e. the set of X{X} corresponding to the same $G_{c}$ is discrete except for a set of $G_{c}$ of measure zero, [19].
@@
Since $V_{o}$ is a dense subset of $\mathbb{C}^{k}$ the dimension of $V_{o}$ is $k$. We define the polynomial matrices $S$ and $T$ such that $T\odot F(X)=S$ where $\odot$ is the Hadamard product of matrices. We let $D$ be the least common multiple of the entries of $T$. Then we define the polynomial matrix $T\odot G_{c}-S$, and the ideal generated by the entries of this matrix and 1tD1-tD. We then compute the Gröbner basis for this ideal using the lexicographic ordering, where the variables in $G_{c}$ should be the smallest variables according to this ordering. The part of the basis containing only the variables in $G_{c}$ defines the smallest variety that contains $V_{c}$, c.f. [20, Ch.3.3]. The dimension of this variety is the same as the dimension of $V_{c}$. Then the algorithm in [21] can be used to compute the dimension.
@@
In case some of the entries of $X$ are known, we just exclude them when defining $V_{o}$ and the variable $X$. When computing the Gröbner basis, one then considers rational functions with coefficients in the field of rational functions of the transfer functions in the known entries of $X$ rather than coefficients in the field $\mathbb{R}$.

### v2

@@ IV Informativity
Here we consider $\Pi$ to just be a function of the free entries of the matrix MM. [^1] When no ambiguity arises we will often just refer to $M$ even if we mean the free entries of $M$. The only property we need regarding these free entries when we investigate generic rank is that they are algebraically independent [^2] Algebraically independent is the formal way of saying that the entries are free. commuting indeterminates over the coefficient field. The coefficient field is the field of rational functions of transfer functions. This is a field with the standard definition of addition and multiplication. We will later on see that the generic rank is equal to the rank for almost all $M$. Since $\Pi$ is a rational function in $z$, it follows that
@@ IV-A Graphical Condition
is a proper affine variety of the set of all $M$. Hence it holds that $\operatorname{rank}_{g}\Pi=\operatorname{rank}\Pi$ for almost all $M$. Then it is straightforward to see that Theorem 2 of that paper can also be generalized. We are therefore able to derive a graph-theoretic characterization of the generic rank of $\Pi$.
@@ IV-B Gröbner Basis Condition
Notice how known entries of $M$ are seamlessly treated since we work over the field of rational functions of transfer functions.
@@ V Identifiability
$F(X)=R_{y}P^{-1}Q_{r}+S_{yr}.$
We let the domain $V_{o}$ of $F$ to be the set of all vectors $X$ of dimension $k$ with elements being rational transfer functions for which the inverse of $P$ exists. This is a dense open subset of the set for which we do not care if the inverse exists or not. We then define the set of admissible $G_{c}$ as $V_{c}=F(V_{o})$. If we take $V_{c}$ to be the co-domain of $F$, it follows that $F$ is a surjective function.
@@
From this definition it can be shown that if a network is generically locally identifiable, then it can be recovered up to a discrete ambiguity, i.e. the set of X{X} corresponding to the same $G_{c}$ is discrete except for a set of $G_{c}$ of measure zero, [19]. Similarly as in this reference we are only referring to one single frequency $z$, and then the norm is well-defined as the complex Euclidian norm. In that paper there is also a discussion about why it is no limitation to just consider one frequency.
@@
Since $V_{o}$ is a dense subset of a set with dimension $k$ it follows that $\dim(V_{o})=k$. We define the polynomial matrices $S$ and $T$ such that $T\odot F(X)=S$ where $\odot$ is the Hadamard product of matrices. We let $D$ be the least common multiple of the entries of $T$. Then we define the polynomial matrix $T\odot G_{c}-S$, and the ideal generated by the entries of this matrix and 1tD1-tD. We then compute the Gröbner basis for this ideal using the lexicographic ordering, where the variables in $G_{c}$ should be the smallest variables according to this ordering. The part of the basis containing only the variables in $G_{c}$ defines the smallest variety that contains $V_{c}$, c.f. [20, Ch.3.3]. The dimension of this variety is the same as the dimension of $V_{c}$. Then the algorithm in [21] can be used to compute the dimension.
@@
In case some of the entries of $X$ are known, we just exclude them when defining $V_{o}$ and the variable $X$.